Suppose the price elasticity of demand for a basket of fruits is -5.0. You are currently charging $42 per basket. If you want to increase the quantity of fruit baskets sold by 20 percent, you should charge a price equal to $_____.  Use midpoint formula.

1
Expert's answer
2022-03-08T12:40:04-0500

Ed=ΔQ%/ΔP%

5=20%/ΔP%

ΔP%=((x-42)/42)×100% =20/-5=-4


(x-42)/42=-4/100=-0.04

x-42=-0.04×42

x-42=-1.68

x=42-1.68=40.32


New price = 40.32 $